# On the Zariski density of rational curves on IHS manifolds

### v2

@@ 4. From $\overline{{\mathbb{Q}}}$ to $\mathbb{C}$ and from weak to strong
The proof is based on [MP12, Theorem 1.1], which states that if we can find a family defined over $\mathbb{K}$, then there are closed points whose Picard rank equals that of the geometric generic fibre. To obtain our claim, it suffices to consider any component $\mathcal{F}$ of the moduli space of polarized IHS manifolds. As these moduli spaces are defined over $\overline{{\mathbb{Q}}}$ (actually over any field of characteristic zero by [And96, Section 2.3.4], therefore we can consider the moduli space ׯ$\mathcal{F}\times_{\overline{{\mathbb{Q}}}}\mathbb{K}$ and any family over a subset of $\mathcal{F}$ whose geometric generic fibre has Picard rank one. Therefore, our claim follows. ∎
@@
Let $X$ be a projective IHS manifold. There exists a constant $d\geq 0$ such that all primitive ample curve classes $[C]\in H_{2}(X,{\mathbb{Z}})$ satisfying $q(C)>d$ have a connected and rational representative $R\in[C]$ such that $R$ rules a prime divisor of class proportional to $[C]^{\vee}$.
@@
Let $\mathcal{M}$ be a component of the moduli space of polarized IHS manifolds and let $(\mathcal{X},\mathcal{H})\xrightarrow{\ \ }\mathcal{M}$ be the universal family over it. Suppose moreover that Conjecture 1.3 and Hypothesis 4.2 hold for all $\overline{{\mathbb{Q}}}$ points of $\mathcal{M}$. Then Conjecture 1.2 holds for all points $t$ of $\mathcal{M}$ for the polarization $\mathcal{H}_{t}$.
@@
Let $t\in\mathcal{M}$. The corresponding polarized $(X_{t},H_{t})$ can be defined over a finite extension of $\overline{{\mathbb{Q}}}$, which implies that we can assume that $(X_{t},H_{t})$ is defined over the function field of an affine variety $B=\textrm{Spec}(A)$ over $\overline{{\mathbb{Q}}}$ ($B$ can of course be a point). We can therefore assume, eventually after shrinking the base, that we have a family $(\mathcal{X},\mathcal{H})\xrightarrow{\ \ }B$ whose geometric generic fibre is $(X_{t},H_{t})$. Let us pick a point $b\in B(\mathbb{Q})$. By hypothesis, we have countably many distinct ruled divisor $\{D_{i}\}_{i\in{\mathbb{N}}}$ inside ${X}_{b}$. Without loss of generality we can suppose that only finitely many of them are not proportional to $H_{b}$, and we can remove these finitely many from the above set. Let us call $C_{i}$ the class of the rational curve ruling $D_{i}$, and let ${h}$ be the relative class of the dual curve to $\mathcal{H}$. For any $i$, there is a $n_{i}$ such that $n_{i}{h}_{b}-C_{i}$ is ample, primitive and has square satisfying hypothesis 4.2. Therefore, we can apply [MP25, Regeneration Principle 1.3] to obtain a regeneration of $D_{i}$ over $B$, which by construction is ruled by a curve of class $n_{i}{h}$. Notice that the $D_{i}$ are all distinct and form a countable set, hence the regenerations we obtain are still countably many. As $(X_{t},H_{t})$ is the generic point of this family, the restrictions of these regenerations to $(X_{t},H_{t})$ are still countably many and Conjecture 1.2 is proven for it. ∎
Notice that in the above proof we might have $A=\mathbb{Q}$, hence the family we consider is a single point. In that case if the manifold $X_{t}$ has Picard rank one, Conjectures 1.2 and 1.3 are equivalent and the latter holds by hypothesis.
@@ 5. Existence from domination
Let $X$ be a projective manifold of OG6 type having a divisor $E$ such that $E^{2}=-2$ and $div(E)=2$. Then $X$ contains infinitely many big and movable ruled divisors.
@@
By [Gro22, Prop. 4.3], we have that $X$ is dominated by a manifold $Y$, which is of $K3^{[3]}$ type and has Picard rank at least $17$, therefore Conjecture 1.3 holds for $Y$ and by Proposition 5.1 the result follows for $X$. ∎
@@
Let $X$ be a projective irreducible symplectic variety contained in tables 4,7 and 8 of [BGMM24], with the exceptions of the 34 items of tables 7 and 8 with rank 7. Then $X$ contains infinitely many big and movable ruled divisors.
@@
By construction, all the irreducible symplectic varieties of [BGMM24] are dominated by manifolds $Y$ of $K3^{[n]}$ or of Kummer type. If the dominating manifold $Y$ has Picard rank at least two, the claim follows by Theorem 1.4 applied to $Y$ and Proposition 5.1. If the action of the group $G$ on $Y$ is non trivial on $H^{2}$, we have an invariant ample class and all coinvariant algebraic classes, therefore the Picard rank is at least two. This holds precisely for all elements in the statement: the rank of the coinvariant lattice is obtained as the difference between the second Betti number and the rank of the invariant part (contained in the above tables), and is positive unless $Y$ is of Kummer type and the rank is 7. ∎

### v3

@@ 4. From $\overline{{\mathbb{Q}}}$ to $\mathbb{C}$ and from weak to strong
The proof is based on [MP12, Theorem 1.1], which states that if we can find a family defined over $\mathbb{K}$, then there are closed points whose Picard rank equals that of the geometric generic fibre. To obtain our claim, it suffices to consider any component $\mathcal{F}$ of the moduli space of polarized IHS manifolds. As these moduli spaces are defined over $\overline{{\mathbb{Q}}}$ (actually over any field of characteristic zero by [And96, Section 2.3.4], therefore we can consider the moduli space $\mathcal{F}\times_{\overline{{\mathbb{Q}}}}\mathbb{K}$ and any family over a subset of $\mathcal{F}$ whose geometric generic fibre has Picard rank one. Therefore, our claim follows. ∎
@@
Let $X$ be a projective IHS manifold. There exists a constant $d\geq 0$ such that all primitive ample curve classes $C]\in H_{2}(X,{\mathbb{Z}})$ satisfying $q(C)>d$ have a connected and rational representative $R\in[C]$ such that $R$ rules a prime divisor of class proportional to $C]^{\vee}$.
@@
Let $\mathcal{M}$ be a component of the moduli space of polarized IHS manifolds and let $(\mathcal{X},\mathcal{H})\xrightarrow{\ \ }\mathcal{M}$ be the universal family over it. Suppose moreover that Conjecture 1.3 (respectively Conjecture 1.2) and Hypothesis 4.2 hold for all $\overline{{\mathbb{Q}}}$ points of $\mathcal{M}$. Then Conjecture 1.3 (respectively 1.2 for the polarization $\mathcal{H}_{t})$ holds for all points $t$ of $\mathcal{M}$.
@@
Let $t\in\mathcal{M}$. If $t$ is defined over $\overline{{\mathbb{Q}}}$, there is nothing to prove. So let us suppose it is not defined over $\overline{{\mathbb{Q}}}$. The corresponding polarized $(X_{t},H_{t})$ can be defined over a finite extension of $\overline{{\mathbb{Q}}}$, which implies that we can assume that $(X_{t},H_{t})$ is defined over the function field of an affine variety $B=\textrm{Spec}(A)$ over $\overline{{\mathbb{Q}}}$ ($B$ can of course be a point). We can therefore assume, eventually after shrinking the base, that we have a family $(\mathcal{X},\mathcal{H})\xrightarrow{\ \ }B$ whose geometric generic fibre is $(X_{t},H_{t})$. Let us pick a point $b\in B(\overline{\mathbb{Q}})$. By hypothesis, we have countably many distinct ruled divisor $\{D_{i}\}_{i\in{\mathbb{N}}}$ inside ${X}_{b}$. Without loss of generality we can suppose that only finitely many of them are not proportional to $H_{b}$, and we can remove these finitely many from the above set. Let us call $C_{i}$ the class of the rational curve ruling $D_{i}$, and let h{h} be the relative class of the dual curve to $\mathcal{H}$. For any $i$, there is a $n_{i}$ such that $n_{i}{h}_{b}-C_{i}$ is ample, primitive and has square satisfying hypothesis 4.2. Therefore, we can apply [MP25, Regeneration Principle 1.3] to obtain a regeneration of $D_{i}$ over $B$, which by construction is ruled by a curve of class $n_{i}{h}$. Notice that the $D_{i}$ are all distinct and form a countable set, hence the regenerations we obtain are still countably many. As $(X_{t},H_{t})$ is the generic point of this family, the restrictions of these regenerations to $(X_{t},H_{t})$ are still countably many and Conjecture 1.2 is proven for it. ∎
Notice that what we actually prove is that Conjecture 1.3 for $\overline{{\mathbb{Q}}}$ points implies Conjecture 1.2 for all other points.
@@ 5. Existence from domination
Let $X$ be a projective manifold of OG6 type having a divisor $E$ such that q(E)=2q(E)=-2 and $div(E)=2$, where $q$ is the Beauville-Bogomolov-Fujiki form. By [Gro22, Prop. 4.3], $X$ is dominated by an IHS manifold $Y$, which is of $K3^{[3]}$ type and has Picard number at least $17$.
@@
Let $X$ be a projective manifold of OG6 type having a divisor $E$ such that q(E)=2q(E)=-2 and $div(E)=2$. If $Y$ has Picard number strictly smaller that $21$, then $X$ contains infinitely many big and movable ruled divisors.
@@
Conjecture 1.3 holds for $Y$ and by Proposition 5.1 the result follows for $X$. ∎
Let $X$ be a projective irreducible symplectic variety contained in tables 4,7 and 8 of [BGMM24]. By construction, each of them is dominated by a IHS manifolds $Y$ of $K3^{[n]}$ or of Kummer type.
@@
Let $X$ be a projective irreducible symplectic variety contained in tables 4,7 and 8 of [BGMM24], with the exceptions of the 34 items of tables 7 and 8 with rank 7. If the dominating variety $Y$ has Picard number strictly smaller than the maximal one, then $X$ contains infinitely many big and movable ruled divisors.
@@
If the dominating manifold $Y$ has Picard rank at least two, the claim follows by Theorem 1.4 applied to $Y$ and Proposition 5.1. If the action of the group $G$ on $Y$ is non trivial on $H^{2}$, we have an invariant ample class and all coinvariant algebraic classes, therefore the Picard rank is at least two. This holds precisely for all elements in the statement: the rank of the coinvariant lattice is obtained as the difference between the second Betti number and the rank of the invariant part (contained in the above tables), and is positive unless $Y$ is of Kummer type and the rank is 7. ∎